OpenStudy (anonymous):

simplify √56x^2 A. 28x B. 2x√14 C. 2x√7 D. 2√14x^2

OpenStudy (*insert name here*):

\[\LARGE \sqrt{56x^2}=\sqrt{4*14*x^2}\] Simplify this:

OpenStudy (anonymous):

a